The half-life for the radioactive decay of C-14 is 5730 years. How long will it take for 25% of the C-14 atoms in a sample of C-14 to decay. If a sample of C-14 initially contains 1.5 mmol of C-14, how many mmols will be left after 2255 years.
Answer #1

K = 0.693/T1/2

     = 0.693/5730   = 0.00012years-1

K = 2.303/t log(a/a-x)

0.00012   = 2.303/t log(100/100-25)

0.00012    =2.303*0.155/t

t                   = 2.303*0.155/0.00012 = 2974.7years

K = 2.303/t log[A0]/[A]

0.00012 = 2.303/2255 log1.5/[A]

log1.5/[A]   = 0.00012*2255/2.303

log1.5/[A] = 0.1175

1.5/[A]    = 100.1175  

1.5/[A]    = 1.3107

[A]      = 1.5/1.3107   = 1.1444m moles>>>> answer